# Uncertainty, Single Crossing Property, and Stochastic Choice Data

### v1

@@ 4.2 Rationalizing information structure and choice rule
$\gamma^{a}(\theta)=\frac{\mu(\theta)q(a|\theta)}{\sum_{\theta^{\prime}\in\Theta}\mu(\theta^{\prime})q(a|\theta^{\prime})}$
@@
$\frac{\gamma^{a^{\prime}}(\theta^{\prime})}{\gamma^{a^{\prime}}(\theta)}=\frac{\mu(\theta^{\prime})}{\mu(\theta)}\frac{q(a^{\prime}|\theta^{\prime})}{q(a^{\prime}|\theta)}\geq\frac{\gamma^{a}(\theta^{\prime})}{\gamma^{a}(\theta)}=\frac{\mu(\theta^{\prime})}{\mu(\theta)}\frac{q(a|\theta^{\prime})}{q(a|\theta)}$
@@
$C(a|\gamma(b))=\begin{cases}\frac{\sum_{\theta^{\prime}\in\Theta}\mu(\theta^{\prime})q(a|\theta^{\prime})}{\sum_{a\in A(b)}\sum_{\theta^{\prime}\in\Theta}\mu(\theta^{\prime})q(a|\theta^{\prime})}&\text{ if }a\in A(b)\\ 0&\text{ if }a\notin A(b)\end{cases}$
@@
The procedure for constructing the information structure is exactly the same as that found in the literature. In particular, we follow the same construction as Caplin and Martin, 2015. An important observation, however, is that when the data is MLR ordered, then so is the rationalizing information structure. That $\langle u,\mathcal{I},C\rangle$ rationalize that data - and hence satisfy conditions 2, 3, and 4 of Definition 4 - follow directly from Theorem 1 of Caplin and Martin, 2015 (see Theorem 4). Since we have shown that $u$ is single crossing and $\mathcal{I}$ is MLR ordered, it follows that the tuple $\langle u,\mathcal{I},C\rangle$ represents an MBEU-DM that rationalizes the data. We have thus shown the following:
@@
Consider the dataset $\{q(\cdot|\theta)\}_{\theta\in\Theta}$. The dataset is MBEU rationalizable if and only if it is MLR ordered.
@@ 5 Discussion
Consider any two actions $a_{k},a_{l}$ and for any $\theta$, denote $A^{\theta}_{k,l}=u^{\theta}_{k}-u^{\theta}_{l}$. As noted above, the proof of sufficiency relies on the main result of Caplin and Martin, 2015, who show that rationalizability by a Bayesian expected utility DM is equivalent to the existence of a solution to a certain system of inequalities (see Appendix A. BEU-Rationalizability). We show that MLR property ensures that a solution exists, which is also single crossing.
@@
where one of two inequalities is strict. Suppose that $u_{k}^{\theta_{N}}-u_{l}^{\theta_{N}}<0$. Then, SCP requires that $u_{k}^{\theta}-u_{l}^{\theta}<0$ for all $\theta$ however, this would imply that numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{k}[u^{\theta}_{k}-u^{\theta}_{l}]\geq 0$. The analyst must then reject the single crossing property. Similarly, if $u_{k}^{\theta_{1}}-u_{l}^{\theta_{1}}>0$, then $u_{k}^{\theta}-u_{l}^{\theta}>0$ for every $\theta$ and therefore the numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{l}[u^{\theta}_{k}-u^{\theta}_{l}]\leq 0$.
However, under uncertainty, almost all of the bite of single crossing property rests upon the information structure being MLR ordered. Since the results here are driven by MCS, it will not be surprising to find that the SCP alone does not impose any strong restrictions. Indeed, one implication of the result of Athey, 2002 is that, if the information structure is not MLR ordered, then there exists a utility function satisfying the SCP such that MCS does not hold. Thus, most of the bite of SCP, in this context, comes when combined with an MLR ordered information structure.
Furthermore, without uncertainty, the only observable implication is monotonicity of chosen actions with respect to the exogenous covariate (e.g. known state), as shown in Lazzati, Quah and Shirai, 2018. As pointed out above, the single crossing property is an ordinal generalization of the property of increasing differences. The latter is intimately related to supermodularity. [^19] A function $f:X\times\Theta\rightarrow\mathbb{R}$ is said to be supermodular if $f(x,\theta)+f(x^{\prime},\theta^{\prime})\leq f((x,\theta)\vee(x^{\prime},\theta^{\prime}))+f((x,\theta)\wedge(x^{\prime},\theta^{\prime}))$, where $X\times\Theta$ is ordered according to the product ordering induced by the orderings over the sets $X$ and $\Theta$. If the functions $x\mapsto f(x,\theta)$ and $\theta\mapsto f(x,\theta)$ are both supermodular for any $\theta$ and $x$ respectively, and if $f$ has increasing differences then $f$ is supermodular. See section 2.6.1 of Topkis, 1998. Interestingly, Chambers and Echenique, 2009 argue that quasi-supermodularity, the ordinal generalization of the notion of supermodularity is indistinguishable from the latter if the underlying domain is finite and the function is assumed to be monotonic. In several settings, therefore, this property is not refutable. The novelty of the results reported here is thus establishing an observational non-equivalence between monotonicity and single crossing property for certain kinds of data.
@@
The proof exhibits that as long as the data is MLR-ordered, one can always find a solution to the system in 13. Therefore, the MLR condition is the only condition required for rationalizing the data by an SCP utility function and MLR information structure. The method of proof here suggests a possible way of approaching problems involving inference in models of information. Existence of a Bayesian DM is equivalent to the existence of a solution to a system of inequalities. Then, suppose an analyst wants to test additional restrictions on these model. The restrictions could either be on the information structure, or the utility function, or both. Whether such restrictions have any implications on stochastic choice data in addition to the NIAS inequalities, is determined by whether these restrictions imply the existence of a solution to the NIAS inequalities, or if they “shrink” the set of solutions. The approach in this paper therefore suggests a method to assess whether a particular restriction is testable.
@@ Appendix A. BEU-Rationalizability
This section briefly presents the result of Caplin and Martin, 2015 since the latter is the result on which our paper builds. A BEU-rationalization of a dataset is $\mathcal{D}$ is a tuple $\langle u,\pi,C\rangle$ such that conditions $2-4$ in Definition 4 hold.
@@
Since $A$ and $\Theta$ are finite, the system of inequalities 13 is a finite system of inequalities. Thus, if there are no solutions to this system, the Bayesian expected utility hypothesis is falsified by observed data. The next section shows that, in binary state case, a single condition on $q\{(\cdot|\theta)\}_{\theta\in\Theta}$ is equivalent to MBEU-rationalization.
@@ Appendix B. Proof for the Binary Case
It can be easily checked that $\langle u,\pi,C\rangle$ is an MBEU-rationalization. Bayes-plausibility follows from construction. Optimality follows from equations 15, 16 and the definition of $(\gamma_{1},\gamma_{2})$. Consistency follows directly by construction of $\pi$. Finally, choosing $(u_{j}^{i})_{i,j\in\{1,2\}}$ such that ${u^{1}_{1}-u^{2}_{1}}>0$ ensures that SCP holds.

### v3

@@ 5.2 Constructing the information structure and the choice rule
$\gamma^{a}(\theta)=\frac{\mu_{0}(\theta)q(a|\theta)}{\sum_{\theta^{\prime}\in\Theta}\mu_{0}(\theta^{\prime})q(a|\theta^{\prime})}$
@@
$\frac{\gamma^{a^{\prime}}(\theta^{\prime})}{\gamma^{a^{\prime}}(\theta)}=\frac{\mu_{0}(\theta^{\prime})}{\mu_{0}(\theta)}\frac{q(a^{\prime}|\theta^{\prime})}{q(a^{\prime}|\theta)}\geq\frac{\gamma^{a}(\theta^{\prime})}{\gamma^{a}(\theta)}=\frac{\mu_{0}(\theta^{\prime})}{\mu_{0}(\theta)}\frac{q(a|\theta^{\prime})}{q(a|\theta)}$
@@
$C(a|\gamma(b))=\begin{cases}\frac{\sum_{\theta^{\prime}\in\Theta}\mu_{0}(\theta^{\prime})q(a|\theta^{\prime})}{\sum_{a\in A(b)}\sum_{\theta^{\prime}\in\Theta}\mu_{0}(\theta^{\prime})q(a|\theta^{\prime})}&\text{ if }a\in A(b)\\ 0&\text{ if }a\notin A(b)\end{cases}$
@@
The procedure for constructing the information structure is exactly as that found in Caplin and Martin, 2015. A key observation, however, is that when the data is MLR-ordered, then so is the rationalizing information structure. That $\langle u,\mathcal{I},C\rangle$ rationalize that data - and hence satisfy conditions 2, 3, and 4 of Definition 4 - follow directly from Theorem 1 of Caplin and Martin, 2015 (see Theorem 5). Since we have shown that $u$ is single crossing and $\mathcal{I}$ is MLR ordered, it follows that the tuple $\langle u,\mathcal{I},C\rangle$ represents an MBEU-DM that rationalizes the data.
We have thus established Theorem 1.
6 Revealed Informedness
Theorem 1 suggests a straightforward, but rather interesting application. Recall that an information structure is said to be Blackwell better than the other if the former guarantees (weakly) highly ex-ante utility for any bounded utility function  [^20] Ex ante utility refers to expected utility before a signal is realized That is, $\sum_{\theta\in S}\sum_{s\in S}\mu_{0}(\theta)\mu(s|\theta)u(a*(s),\theta))$. This is often referred to as the value of the information structure.. Indeed, since Blackwell order requires an information structure to give higher ex-ante value for every utility function, it is, in general, a severely partial order. By restricting attention to smaller classes of utility functions, more complete orders over information structures can be obtained.
For single crossing utilities, we typically consider the Lehmann order (Lehmann, 1988), which is weaker than the Blackwell order. In particular, if an information structure $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in Blackwell order, then $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in Lehmann order, but the converse may ot hold. In particular, $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in the Lehmann order if and only if for every single-crossing utility function ex-ante utility is higher under $\mathcal{I}_{1}$ than under $\mathcal{I}_{2}$ (see for example, Persico, 2000, Jewitt, 2006). Below we give a formal definition of Lehmann order.
Definition 5.
Let $\mathcal{I}_{i}\equiv\{\mu^{i}(\cdot|\theta)\}_{\theta\in\Theta}\subset\Delta(S^{i}),i=1,2$ be MLR-ordered information structures. Let $\{F^{i}_{\theta}(\cdot)\}_{\theta\in\Theta},i=1,2$ be the corresponding distribution functions. Then, $\mathcal{I}_{1}$ is said to be Lehmann higher, or more accurate, than $\mathcal{I}_{2}$ if there exists an function $h:S^{2}\times\Theta\rightarrow S^{1}$ which is increasing in $\theta$ and, such that, for every $\theta,s^{1}\in S^{1}$, [^21] While this definition is appropriate only for continuous distributions $F^{i}$, Lehmann, 1988 shows that it is without loss of generality to assume that $F^{i}$ are continuous. That is, if $F^{1},F^{2}$ are discontinuous, there exist continuous distributions $F^{i^{*}}$ which are informationally equivalent to $F^{i}$.
$F^{1}_{\theta}(h(s^{2},\theta))=F^{2}_{\theta}(s^{2}).$
Suppose $\mathcal{I}_{1}=\{\mu^{1}(\cdot|s)\}_{s\in S^{1}},\mathcal{I}_{2}\equiv\{\mu^{2}(\cdot|s)\}_{s\in S^{2}}$ are MLR ordered, and let $\nu^{1},\nu^{2}$ be the corresponding marginal distributions over signals. [^22] So, $\nu^{i}(s)=\int_{\theta\in\Theta}\mu^{i}(s|\theta)\mu_{0}(\theta)$ for each $s\in S^{i}$. $\mathcal{I}_{1}$ is Lehmann higher than $\mathcal{I}_{2}$ if and only if, for every single crossing utility function $u(a,\theta)$, [^23] See Lehmann, 1988, theorem 5.1.
$\int_{s\in S^{1}}\int_{\theta\in\Theta}u(a^{*}(s),\theta)d\mu^{1}(\theta|s)d\nu^{1}(s)\geq\int_{s\in S^{2}}\int_{\theta\in\Theta}u(a^{*}(s),\theta)d\mu^{2}(\theta|s)d\nu^{2}(s).$
Our main result gives the conditions under which rationalizing posteriors and single crossing utility functions may be constructed from the observed state conditional stochastic choice data. Suppose that two distinct datasets are MBEU-rationalizable. If the analyst can somehow conclude that one rationalizing information structure, say $\mathcal{I}_{1}$ is Lehmann higher than the other, $\mathcal{I}_{2}$, then, the analyst can infer that any MBEU-DM rationalizing the former has better information than the one rationalizing the latter.
Definition 6.
Fix some class $\mathcal{U}$ of utility functions, and consider two datasets $\mathcal{D}^{1}=\{q^{1}(\cdot|\theta)\}$, $\mathcal{D}^{2}=\{q^{2}(\cdot|\theta)\}$. We say that the DM corresponding to $\mathcal{D}^{1}$ is $\mathcal{U}$-revealed more informed than the DM corresponding to $\mathcal{D}^{2}$ if, there exist information structures $\mathcal{I}_{1},\mathcal{I}_{2}$ such that, for any utility functions $u_{1},u_{2}\in\mathcal{U}$, if $(\mathcal{I}_{1},u_{1}),(\mathcal{I}_{2},u_{2})$ are BEU-rationalizations of $\mathcal{D}^{1},\mathcal{D}^{2}$, respectively, then,
$\mathbb{E}_{\mathcal{I}_{1}}u_{1}\geq\mathbb{E}_{\mathcal{I}_{2}}u_{1}$
$\mathbb{E}_{\mathcal{I}_{1}}u_{2}\geq\mathbb{E}_{\mathcal{I}_{2}}u_{2}$
Clearly, if $\mathcal{U}$ is the family of all single crossing utility functions, the MBEU-DM corresponding to $\{q^{1}(\cdot|\theta)\}$ is revealed more informed than the MBEU-DM corresponding to $\{q^{2}(\cdot|\theta)\}$ if and only if $\mathcal{I}_{1}$ is Lehmann higher than $\mathcal{I}_{2}$. Then, we have the following result:
@@ 6 Revealed Informedness
Let $\mathcal{D}^{1}=\{q^{1}(\cdot|\theta)\},\mathcal{D}^{2}=\{q^{2}(\cdot|\theta)\}$ be MLR ordered and such that $\mathcal{D}^{1}$ is Lehmann higher than $\mathcal{D}^{2}$. Then, the MBEU-DM corresponding to $\mathcal{D}^{1}$ is revealed more informed than the MBEU-DM corresponding to $\mathcal{D}^{2}$.
Proof.
See appendix. ∎
While the proof is entirely trivial, we contend that this result is nonetheless interesting. To underscore the argument, it is instructive to compare our rationalizability result with that of Caplin and Martin (2015). Indeed, if one is interested in BEU-rationalizability alone, and the analyst has datasets $\mathcal{D}^{1},\mathcal{D}^{2}$ such that $(i)$ both $\mathcal{D}^{1},\mathcal{D}^{2}$ are BEU-rationalizable and $(ii)$ $\mathcal{D}^{1}$ Blackwell dominates $\mathcal{D}^{2}$, then the analyst can conclude that the BEU-DM corresponding to $\mathcal{D}^{1}$ is more informed than the BEU-DM corresponding to $\mathcal{D}^{2}$. But, as pointed out above, it is not known whether, in general, a solution to the NIAS inequalities exists. Secondly, Blackwell order is quite incomplete. On the other hand, the definition of Lehmann order requires the datasets to be MLR ordered - which guarantees the existence of an MBEU-DM.
@@ 7 Discussion
Consider any two actions $a_{k},a_{l}$ and for any $\theta$, denote $A^{\theta}_{k,l}=u^{\theta}_{k}-u^{\theta}_{l}$. As noted above, the proof of sufficiency relies on the main result of Caplin and Martin, 2015, who show that rationalizability with a Bayesian expected utility DM is equivalent to the existence of a solution to a certain system of inequalities (see Appendix A. BEU-Rationalizability). We show that MLR property ensures that a solution exists, which is also single crossing.
@@
where one of two inequalities is strict. Suppose that $u_{k}^{\theta_{N}}-u_{l}^{\theta_{N}}<0$. Then, SCP requires that $u_{k}^{\theta}-u_{l}^{\theta}<0$ for all $\theta$ - however, this would imply that numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{k}[u^{\theta}_{k}-u^{\theta}_{l}]\geq 0$. The analyst must then reject the single crossing property. Similarly, if $u_{k}^{\theta_{1}}-u_{l}^{\theta_{1}}>0$, then $u_{k}^{\theta}-u_{l}^{\theta}>0$ for every $\theta$ and therefore the numbers $u^{\theta}_{k},u^{\theta}_{l}$ cannot be a solution to the inequality $\sum_{\theta}q^{\theta}_{l}[u^{\theta}_{k}-u^{\theta}_{l}]\leq 0$. However, under uncertainty, almost all of the bite of single crossing property rests upon the information structure being MLR ordered. Indeed, one implication of the result of Athey, 2002 is that, if the information structure is not MLR ordered, then there exists a utility function satisfying the SCP such that MCS does not hold. Thus, under uncertainty, most of the bite of SCP comes when combined with an MLR ordered information structures.
Furthermore, without uncertainty, the only observable implication is monotonicity of chosen actions with respect to the exogenous covariate (e.g. known state), as shown in Lazzati, Quah and Shirai, 2018. As pointed out above, the single crossing property is an ordinal generalization of the property of increasing differences. The latter is intimately related to supermodularity. [^24] A function $f:X\times\Theta\rightarrow\mathbb{R}$ is said to be supermodular if $f(x,\theta)+f(x^{\prime},\theta^{\prime})\leq f((x,\theta)\vee(x^{\prime},\theta^{\prime}))+f((x,\theta)\wedge(x^{\prime},\theta^{\prime}))$, where $X\times\Theta$ is ordered according to the product ordering induced by the orderings over the sets $X$ and $\Theta$. If the functions $x\mapsto f(x,\theta)$ and $\theta\mapsto f(x,\theta)$ are both supermodular for any $\theta$ and $x$ respectively, and if $f$ has increasing differences then $f$ is supermodular. See section 2.6.1 of Topkis, 1998. Interestingly, Chambers and Echenique, 2009 argue that quasi-supermodularity, the ordinal generalization of the notion of supermodularity is indistinguishable from the latter if the underlying domain is finite and the function is assumed to be monotonic. In several settings, therefore, this property is not refutable. A novelty of the results reported here is thus establishing an observational non-equivalence between monotonicity and single crossing property for certain kinds of data.
@@
The proof exhibits that as long as the data is MLR-ordered, one can always find a solution to the system in 14. Therefore, the MLR condition is the only condition required for rationalizing the data by an SCP utility function and MLR information structure. The method of proof here suggests a possible way of approaching problems involving inference in models of information. Existence of a Bayesian DM is equivalent to the existence of a solution to a system of inequalities. Then, suppose an analyst wants to test additional restrictions on this model. The restrictions could either be on the information structure, or the utility function, or both. Whether such restrictions have any implications on stochastic choice data in addition to the NIAS inequalities, is determined by whether these restrictions imply the existence of a solution to the NIAS inequalities, or if they “shrink” the set of solutions. The approach in this paper therefore suggests a method to assess whether a particular restriction is testable.
@@ Appendix A. BEU-Rationalizability
This section briefly presents the result of Caplin and Martin, 2015 since the latter is the result on which our paper builds. A BEU-rationalization of a dataset is $\mathcal{D}$ is a tuple $\langle u,\pi,C\rangle$ such that conditions 242-4 in Definition 4 hold.
@@
Since $A$ and $\Theta$ are finite, the system of inequalities 14 is a finite system of inequalities. Thus, if there are no solutions to this system, the Bayesian expected utility hypothesis is falsified by observed data. The next section shows that, in binary state case, a single condition on $q\{(\cdot|\theta)\}_{\theta\in\Theta}$ is equivalent to MBEU-rationalization.
@@ Appendix B. Proof for the Binary Case
It can be easily checked that $\langle u,\pi,C\rangle$ is an MBEU-rationalization. Bayes-plausibility follows from construction. Optimality follows from equations 16, 17 and the definition of $(\gamma_{1},\gamma_{2})$. Consistency follows directly by construction of $\pi$. Finally, choosing $(u_{j}^{i})_{i,j\in\{1,2\}}$ such that ${u^{1}_{1}-u^{2}_{1}}>0$ ensures that SCP holds.
@@
Appendix C. Proof of Theorem 4
Proof.
Consider the datasets $\mathcal{D}^{i}\{q^{i}(\cdot|\theta)\}_{\theta\in\Theta}$, $i=1,2$, and let $F^{i}_{\theta}$ be the cumulative distribution corresponding to $q^{i}(\cdot|\theta)$. From Lehmann, we know that it is without loss of generality to assume that $F^{i}_{\theta}$ is continuous for each $\theta$. So, $F^{i}_{\theta}(a)=\int_{a^{\prime}<a}q^{i}(a^{\prime}|\theta)da^{\prime}$. Recall that, the constructed posterior rationalizing the dataset was defined as $\gamma^{a}_{i}(\theta)=\frac{q^{i}(a|\theta)\mu_{0}(\theta)}{\sum_{\theta^{\prime}\in\Theta}q^{i}(a|\theta^{\prime})\mu_{0}(\theta)}=\frac{q^{i}(a|\theta)\mu_{0}(\theta)}{\nu^{i}(a)}$ - where $\nu^{i}$ is the marginal distribution over actions corresponding to $\mathcal{D}^{i}$. But then, the probability of obtaining a signal $a$ when the state is $\theta$ is simply $\gamma_{i}(\theta|a)=\frac{\gamma^{a}_{i}(\theta)\nu^{i}(a)}{\mu_{0}(\theta)}=q^{i}(a|\theta)$. Therefore, the cumulative distribution of $\gamma_{i}(a|\theta)$ - the state conditional distributions over signal realizations - is simply $F^{i}_{\theta}$, and thus the rationalizing information structures are Lehmann ordered. That is, the MBEU-DM corresponding to $\mathcal{D}^{1}$ is revealed more informed than the MBEU-DM corresponding to $\mathcal{D}^{2}$. ∎